Your monthly mortgage payment (principal plus interest) is $1,750 . If you have a 30-year loan with a fixed interest rate of 0.3% per month, how much did you borrow from the bank to purchase your house (rounded to the nearest whole dollar)? (Do not enter a dollar sign $ with your answer.)
Answer #1

total interest payable ( denoting it with I)

I= P*R*T /100 where P is principal

R is interest rate charged, T is time period in which loan has to be repaid.

suppose borrowed amount be P

R= 0.3%, T=30 yrs

I = P * 0.3 *30 *12/100 = 108 P/100 = total interest we'll pay.

total amount will be = P+ 108 P/100

we've per month payment of 1750

making total amount into what has to be paid in one month = P (1+1.08)/ 30*12

equating this amount to 1750.

P = 1750 *30*12/2.08 = 302884.615 = borrowed amount